### v1

# thanks: These authors contributed equally to this work

Abstract
The original proposal of quantum key distribution (QKD) was based on ideal single photon sources, which 40 years later, are still challenging to develop. Therefore, the development of decoy state protocols using weak coherent states (WCS) from lasers, set the frontier in terms of secure key rates. Here, we propose and experimentally emulate two simple-to-implement protocols that allow practical, far from ideal sub-Poissonian photon sources to outperform state-of-the-art WCS. By engineering the photon statistics of a quantum dot’s biexciton-exciton cascade, we show that either a truncated decoy state protocol or a heralded purification protocol can be employed to achieve a significantly increased performance in terms of the maximal allowed channel loss for secure key creation, which can exceed that of WCS by more than 3dB. We then show that our recently demonstrated room temperature single photon sources, based on giant colloidal quantum dots coupled to nano-antennas, are already well within the optimal performance range. These protocols can be utilized efficiently on a host of various sub-Poissonian quantum emitters having controllable photon statistics, offering a practical approach to QKD without the hindering requirements on the single photon purity of the photon source.
I Introduction
Quantum key distribution (QKD) stands as a prominent candidate for the post-quantum-computer secure optical communication, enabling a secure key-establishing protocol between two parties employing the quantum nature of light [1, 2, 3]. In theory, encoding the information for QKD onto pure single photon states, is ultimately secured. In practice, it is limited by imperfections of the transmitter and the receiver, making security proofs a challenging task [4, 5], thus reducing the maximal secure key rate (SKR) and the maximal losses in the communication channel over which SKR is achievable [6, 7].
One of the main practical limitations of QKD systems is the lack of an ideal source for single photons (SPS), required to prevent eavesdropping attacks such as photon number splitting attack (PNS) [8]. The SKR and maximal allowed channel loss (MCL) of simple QKD protocols decrease dramatically with the increase of two-or-more photon events being transmitted from the source [9, 10]. This leads to very stringent requirements on both the single photon purity and the photon emission rate of any practical SPS [11].
Despite many years of exploration and engineering [12, 13], a simple, stable SPS system that has the required aforementioned properties and can be employed in real-life QKD applications is still an outstanding challenge, which led researchers to suggest and utilize protocols employing attenuated lasers emitting weak coherent states (WCS) [14]. As WCS has a Poissonian distribution of photons, and thus a two-or-more photons probability is never vanishing, advanced protocols such as decoy state protocols have been developed to enhance both SKR and MCL [11, 15, 16, 17, 18].
Decoy state protocols introduce variations in the distribution of the emitted photons to better characterize the channel and obtain a tighter bound on the SKR [19, 20, 21, 22], allowing detection of PNS attacks [23, 24]. Yet, even with decoy state protocols, WCS are highly sensitive [25, 26] to the high probability of the vacuum state, due to the dependence between the different photon number emission probabilities, which overall limits the possible SKR [9]. The above solutions and inherent limitations of WCS for QKD set clear opportunities for improving the security and efficiency of QKD applications using realistic SPS systems [27].
While efforts to develop a nearly ideal SPS are still ongoing by many groups [28, 29, 30, 31, 32], here we introduce a radically different, more realistic approach. By utilizing two new simple QKD protocols on compact, imperfect SPS sources, we significantly enhance the QKD performance, in terms of the MCL, as compared to even ideal, infinite decoy state protocols using WCS.
We base our protocols on our recent demonstrations of room temperature operating SPS devices based on a giant colloidal nanocrystal quantum dot (gCQD) coupled to a hybrid nanoantenna and plasmonic nanocone [33, 34, 35, 36, 37]. These SPS devices displayed a highly enhanced rate of photon emission approaching the GHz range, together with near-unity collection efficiencies ($>90\%$) of the emitted photons. The photon emission from such gCQDs is based on the bi-exciton - exciton (BX-X) emission cascade under optical excitation [38], and has been shown to be well described by a truncated photon-number Fock space with $N=0,1$, and $2$ photons, with a negligible probability of $N>2$ photons [39].
The two protocols are both based on our ability to vary and control the probabilities of emitting the different Fock states, $\{P_{0},P_{1},P_{2}\}$ over a wide range of values, by simply varying the optical excitation power. The first protocol we implement is a simple and realistic decoy state protocol for such a source. We show that with this new protocol, we can significantly exceed, by more than $3$ dB, the MCL value of an ideal decoy state protocol using a WCS source. The second protocol uses an alternative approach based on heralded photon purification under saturated excitation. We show that a similar performance enhancement can also be achieved. Finally we experimentally verify that the probability for $N>2$ ($P_{>2}$) is negligibly small, justifying our protocols’ assumptions, and demonstrating that even existing SPS which are far from being ideal, can outperform the current state-of-the-art protocols based on WCS sources.
Figure 1: (a) Schematic of the BX-X cascade in an optically excited gCQD, using a non-resonant pulsed excitation with a normalized intensity $S=\frac{I}{I_{S}}$. Inset is a sketch of a typical gCQD, consisting of a CdSe core and a CdS shell. (b) Normalized emission counts (blue open circles) as a function of $S$, displaying a saturation behaviour. The dashed black line is a fit to a model, detailed in Section II. The orange circles show the experimentally extracted $g^{(2)}(0)$ for the different powers. (c) Photon probabilities as a function of $S$. The dashed black lines mark the intensities chosen for the signal $S_{2}$, and the two decoy states $S_{0}$, $S_{1}$, for the decoy state on a truncated basis (DTB) protocol.
II Optical control of photon statistics from the BX-X emission cascade in a gCQD
Fig. 1a presents a schematic sketch of the BX-X cascade in a gCQD (see Supp. Sec. I) following a non-resonant pulse excitation with power $I$. Modelling this process, a BX state |$\ket{XX}$ is excited by the absorption of two pump photons. This process happens with a probability:
$P_{XX}=\frac{\left(\alpha I\right)^{2}}{1+\alpha I+\left(\alpha I\right)^{2}}.$ (1)
The probability of excitation of only the X state, |$\ket{X}$, by the absorption of only one pump photon, is given by:
$P_{X}=\frac{\alpha I}{1+\alpha I+\left(\alpha I\right)^{2}}.$ (2)
Here $\alpha$ is a constant related to the absorption cross-section of the gCQD and we assume that any larger excitonic complexes are negligible [39], as will be verified later in this paper.
Following an optical excitation, the biexciton state |$\ket{XX}$ decays into a single exciton state |$\ket{X}$ either radiatively, by emitting a single photon with a probability $QY_{XX}$, representing the BX quantum yield, or non-radiatively with a probability $1-QY_{XX}$. The exciton state |$\ket{X}$ can then recombine radiatively (non-radiatively) with probabilities $QY_{X}$, ($1-QY_{X}$), to the ground state such that the probability to emit zero, one or two photons, defined as $\{P_{0},\,P_{1},\,P_{2}\}$ respectively, is:
$P_{1}$ $=P_{XX}\left(QY_{X}+QY_{XX}-2QY_{X}QY_{XX}\right)$ (3)
$+P_{X}QY_{X}$
$P_{2}$ $=P_{XX}QY_{X}QY_{XX}$ (4)
$P_{0}$ $=1-P_{1}-P_{2}$ (5)
Since $P_{XX},P_{X}$ depend on the excitation power, the probabilities to emit two, one or zero photons following excitation become excitation power dependent, allowing an external control over the emitted photon statistics as is required for decoy states protocols. Such a saturable behavior of the excitonic population allows us to define a relative excitation power $S=I/I_{S}$, where $I_{S}$ is the saturation power, for which the detected count rate (which is linearly proportional to the average emission photon number) reaches $90\%$ of the maximum value.
In Fig. 1(b), we experimentally demonstrate photon statistics control by varying the excitation power from a single bare gCQD. As $S$ increases, the collected photon emission intensity increases sublinearly and saturates at $S\gtrsim 1$. A good fit of the theoretical model, based on the absorption probabilities given in Eqs. 1, 2 and on the subsequent emission probabilities $QY_{X},QY_{XX}$ is shown (see Supp. Sec. II). This fit confirms the basic assumptions given above for the BX-X cascade process.
At the same time, the measured second-order correlation function, $g^{(2)}(0)$, also increases in a similar manner. This is understood in the following way: at low powers, most emission events are empty as $P_{0}$ is the dominant term. As $S$ increases, $P_{1}$ increases linearly (Eq. 2), while $P_{2}$ increases quadratically and thus is negligible at $S\ll 1$. Since $g^{(2)}(0)$ is an increasing function of $P_{2}/(\sum_{i=0,1,2}{P_{i}})$, a saturable behavior of $g^{(2)}(0)$ with increasing $S$ is also expected for $S\rightarrow 1$.
Importantly, as we also show experimentally, multi-excitonic emissions from the gCQD can be neglected due to their low emission probability [39], thus $P_{N>2}=0$.
By measuring both $g^{(2)}(0)$ and the count-rate $C$, we can extract the probability distribution precisely, as is detailed in Supp. Sec. III.
In Fig. 1(c), we present the extracted photon emission probabilities $\{P_{0},P_{1},P_{2}\}$. As expected, the probability of emitting one or two photons increases with excitation power, but at a different rate, allowing optical control of the emitted photon statistics with only varying $S$.
Figure 2: Concept of two protocols based on a high rate, high collection efficiency SPS with imperfect purity: (a) Experimental setup and sketch of the BB84+DTB protocol, with Alice adjusting the excitation power on the SPS between $S_{0}=0$, $S_{1}$ and $S_{2}$ to control the emitted photon statistics for the 2-decoy protocol, followed by collection optics with efficiency $\eta_{C}$ and a standard BB84 encoding unit. (b) Experimental setup and sketch of the BB84+HPP photon purification protocol. A single excitation power $S_{3}>1$, near saturation, is used to excite the SPS followed by collection optics with efficiency $\eta_{C}$ and a beam-splitter (BS) with a high efficiency single photon detector with detection efficiency $\eta_{D}$. Here only same pulse detections at Alice and Bob are used for the secure key.
III QKD Protocols outperforming WCS
After establishing the control of the different photon number probabilities, we now utilize this capability for two MCL-enhancing protocols, the first is a decoy-state protocol and the second is a heralded purification protocol.
III.1 Decoy State on a Truncated Fock Basis - DTB
In Fig. 2(a), we present the conceptual experimental setup of the decoy state protocol. Here an SPS with a near unity photon collection efficiency, $\eta_{c}\simeq 1$ [36] (titled here SPS1), is used for implementing a BB84 protocol, but where Alice controls the excitation powers, $S_{0}=0$, $S_{1}$ and $S_{2}$. This in turn modifies the photon number statistics, thus allowing for the implementation of decoy and signal states [11, 15].
To establish the DTB protocol, we define the gain of the signal (decoyi), $Q_{s}$ ($Q_{d,i}$), as the fraction of encoded photon pulses sent by Alice and detected by Bob. We also define the error rate of the signal (decoyi), $E_{s}$ ($E_{d,i}$), as the fraction of detected encoded photons that had errors. The signal pulses correspond to pulses with excitation $S_{2}$, while the decoy pulses are those excited with $S_{1},S_{0}$. We show in the following that due to the truncated photon number basis of our SPS, only two different decoy states are enough for exact analysis in the framework of the decoy state protocol [11].
Within the DTB protocol, Alice can randomly choose to replace the signal pulse with one of the two decoy states with different photon number distributions. These decoy states are not used to create the secure key but to improve the information about the channel and discover any possible eavesdroppers [15]. In the post-processing stage, Alice and Bob can verify the fraction of detected events and the errors for both signal pulses and decoy pulses, thus obtaining the gain and error rates $\{Q_{s},Q_{d,i},E_{s},E_{d,i}\}$.
On the other hand, due to the uncertainty of the photon-number in each pulse, the gain and error rate for a specific n-photon state, $\{Q_{n},e_{n}\}$, cannot be measured directly and have to be estimated or calculated. Out of this set, the single photon gain and error rate, $Q_{1},\,e_{1}$ which are the probability of a detected event to originate from a single photon pulse and the error rate for single photon pulses respectively are of particular importance. This is since, considering a possible PNS attack by Eve, only information encoded on the single photon pulses is secured.
Crucially, in our three-intensity DTB protocol, these parameters can be solved exactly since the probability distribution of the signal and decoy states, $\{P_{n}^{s}\},\{P_{n}^{d,i}\}$ is known, e.g. Fig. 1c. To relate the measured values $\{Q_{s},Q_{d,i},E_{s},E_{d,i}\}$ to the unknown set $\{Q_{n},e_{n}\}$, it is useful to define the n-photon yield, $Y_{n}$, as the conditional probability of a detection event given that Alice sent a n-photon state, and therefore $Q_{n}=Y_{n}P_{n}$. With these definitions, the following set of equations can be written [15]:
$Q_{s}$ $=\sum_{n=0}^{2}P_{n}^{s}Y_{n}\qquad E_{s}Q_{s}=\sum_{n=0}^{2}P_{n}^{s}Y_{n}e_{n}$ (6)
$Q_{d,i}$ $=\sum_{n=0}^{2}P_{n}^{d,i}Y_{n}\qquad E_{d,i}Q_{d,i}=\sum_{n=0}^{2}P_{n}^{d,i}Y_{n}e_{n}$ (7)
Importantly, unlike the decoy state protocol implemented for WCS where $n\rightarrow\infty$, and therefore require infinite set of decoy states for an exact solution [11], here $n=0,1,2$ only. Thus, our equations have only six unknowns: $\{Y_{0},\,e_{0},\,Y_{1},\,e_{1},\,Y_{2},\,e_{2}\}$ suggesting that two decoy states are enough for an exact solution to these equations. This makes the DTB implementation particularly viable, whereas for WCS using two decoy states would only give a bounded approximation for the yield and errors [15].
The solution of these equations can then be used to estimate the minimum SKR after privacy amplification and error correction [11]:
$SKR\geq q\{-Q_{s}f(E_{s})H_{2}(E_{s})+Q_{1}\left[1-H_{2}(e_{1})\right]\}$ (8)
where $q=0.5$ for the BB84 protocol [1], $H_{2}$ is the binary Shannon information function [11] and $f(E_{s})$ is the error correction efficiency (taken to be 1.22 [40]).
Assuming $\eta_{C}\simeq 1$ for gCQD on a nanoantenna [36] and using the probability values- $P_{0},P_{1},P_{2}$, we can numerically calculate the SKR of such a source for different channel losses. For this, we use the channel model connecting the losses, yields, gains and the error-rates shown in [11] and in Supp. Sec. IV with the parameters values given in [41, 42], specifically, dark count probability of $P_{\text{dark}}=1.7\cdot{10}^{-6}$, probability to reach the wrong detector of $e_{d}=3.3\%$, Bob’s detection efficiency of $\eta_{\text{Bob}}=4.5\%$ and vacuum error-rate of $e_{0}=0.5$.
Figure 3: QKD analysis with imperfect single photon sources. (a) Secure key rate of BB84 protocol as a function of channel loss for a perfect SPS (blue line), WCS with infinite decoy states and optimized intensities (WCS+Decoy, red), SPS1 with the decoy state on truncated basis protocol (SPS1+DTB, purple) and SPS2 with the heralded purification protocol (SPS2+HPP, black). The method and parameters of the calculations are found in Sec. III.1 and III.2. (b) The relative gain of the MCL, $\gamma$, as a function of $P_{1},P_{2}$ showing two advantageous regimes for SPS with either DTB or HPP separated by black dashed lines with their respective conditions. We also present our sources, SPS1 in the purification regime (black dot), SPS2 in the decoy state regime (purple square) and two, previously analyzed, non-classical sources [42, 43] marked with the orange square/dot.
In Fig. 3(a) we show the calculated SKR under different channel losses for a standard WCS with decoy states (red) in the ideal, asymptotic case (infinite number of decoy states) with optimized intensity, compared to an SPS based on a gCQD coupled to a nanoantenna [36] (purple) using the above DTB protocol, with the realistically obtainable probabilities for such a device: $P_{1}^{s}=0.529$, $P_{2}^{s}=0.112$ [36] for the signal state. A clear improvement in the SKR and the MCL of our imperfect SPS over WCS is seen, with over $3$ dB MCL enhancement. We also show a comparison to a perfect single photon source ($P_{1}=1$) in black. Surprisingly, applying our truncated decoy-state protocol yields performance not far worse than a perfect SPS, even though our SPS is far from being ideal.
III.2 Heralded Purification Protocol- HPP
In Fig. 2(b), we present the experimental configuration of the heralded purification protocol (BB84+HPP). In this scheme, we consider a SPS consisting for example a gCQD on a hybrid nanocone-antenna device (titled here SPS2). Such a device showed both $\eta_{C}\simeq 1$ together with high attainable values of $P_{2}$, resulting from a large Purcell factor induced by the nanocone which significantly enhances $QY_{XX}$ [34, 44]. Alice operates at a single excitation power $S_{3}>1$, deep in the saturation regime, to excite the BX state with very high probability, thus maximizing $P_{2}$, which is now only limited by $QY_{XX}$ and $QY_{X}$ [34]. A beam-splitter (BS) and a single photon detector (SPD) are added as a purification stage after the SPS in Alice’s system, and before the standard BB84 encoding unit. As $P_{N>2}$ are negligible (see Supp. Sec. III), a real photon detection event in Alice’s detector sets $\tilde{P}_{2}=0$ for that pulse, where $\tilde{P}_{n}$ is the effective photon number probability after the purification stage. In the sifting step of the QKD protocol, only events with same-pulse detections by Alice and Bob are considered for the secure key, thus eliminating all multiphoton events at a cost of lower signal rates.
Figure 4: Experimental results for emulating the DTB protocol. (a) Controlling the photon emission counts by the excitation laser intensity (405 nm, 2 MHz rep. rate, 2 ns pulse duration). Red regions mark $S_{0}$, green regions mark $S_{1}$ and blue regions mark $S_{2}$ appearing in Fig. 1(c). (b) $g^{(2)}(0)$ for the two different intensities, $S_{1}$, $S_{2}$. (c) Calculation of $\gamma$ values as a function of $\eta_{C}$. The blue box represents the $\eta_{C}$ range demonstrated in our devices [34, 36], all with $\gamma>2$ dB.
Figure 5: Experimental results for emulating the HPP protocol. Second-order correlation measurements of the bare gCQD SPS without (a) and with (b) HPP post-processing scheme, resulting in near-zero $g^{(2)}(0)$. The error is mainly due to the SPD dark noise. (c) Calculation showing the optimal transmission $T$, of Alice’s BS, required for maximizing $\gamma$ as a function of $P_{2}$ of the SPS, for two values of $P_{DC}$, corresponding to 100 dark counts per second of the SPD, and SPS excitation repetition rates of $2$ (blue) and $500$ (red) MHz. (d) Calculation of $\gamma$ as a function of $\eta_{C}$ with a fixed $\eta_{D}$ (blue line) and as a function of $\eta_{D}$ with a fixed $\eta_{C}$ (red line). The blue box represents the range demonstrated in our devices [34, 36].
Given a SPS having $\{P_{0},\,P_{1},\,P_{2}\}$, the effective distribution sent to Bob with HPP depends on the reflectance and transmission ($R$ and $T$) of the BS and the detection efficiency of Alice’s detector ($\eta_{D}$), as well as the probability of a dark count at Alice’s detector ($P_{DC}$):
$\tilde{P}_{1}$ $=2P_{2}RT(\eta_{D}+P_{DC})+TP_{1}P_{DC}$ (9)
$\tilde{P}_{2}$ $=T^{2}P_{2}P_{DC}$ (10)
$\tilde{P}_{0}$ $=1-\tilde{P}_{1}-\tilde{P}_{2}$ (11)
Given this new distribution, we can again follow the well established method to estimate the SKR of a BB84 protocol (but now implemented with HPP) [11, 45]. The black line in Fig. 3(a) shows a realistic calculation of the SKR using actual parameters of SPS2 measured in Ref. [34]. A BS reflectance of $50\%$ was chosen. We also used a SPD with $\eta_{D}=0.9$ and $P_{DC}=2\cdot{10}^{-7}$ corresponding to 100 dark counts per second for a $500$ MHz signal rate which are both commonly attainable with current technology. Again, as is seen in the figure, the new BB84+HPP allows for a higher MCL ($\sim 1$ dB) compared to WCS with infinite decoy, due to the extremely low $\tilde{P}_{2}$. Remarkably, again this SKR enhancement can be achieved with SPS far from being ideal having a very low single photon purity.
III.3 Performance Analysis
Next, we analyze the expected performance of realistic SPS-based BB84-QKD using the above protocols, in comparison to that of WCS with ideal decoy protocols. In particular, we compare the expected performance of our existing room-temperature high brightness, high collection efficiency SPS [34, 36]. We define the relative gain in the MCL as $\gamma=(MCL/MCL_{WCS})$, and use this parameter to evaluate the relative performance gain.
Fig. 3(b) shows a colormap of the calculated $\gamma$ for different values of $P_{1},P_{2}$ of the SPS (see Supp. Sec. IV). For high $P_{1}$ and sufficiently low $P_{2}$, namely for $P_{1}>1.125P_{2}+0.1927$, there is a region (marked as ”SPS+DTB”) where $\gamma>0\,\text{dB}$, indicating that the use of an SPS with BB84+DTB protocol is advantageous over BB84 including decoy states with WCS (WCS+D). In this region, we highlight in orange dots two known non-classical sources [42, 43] along with our device (SPS1) consisting of a gCQD coupled to a metal-dielectric Bragg nanoantenna [36] (purple), demonstrating that already existing sources, when combined with the DTB protocol, can outperform even ideal WCS protocols. Notably, the use of DTB allows for SPS with a higher probability of two-photon events and shows that bright devices (with fewer vacuum events) can be used for QKD even with single photon purities as low as $\sim 65\%$ and $g^{(2)}(0)$ values as high as $\sim 0.6$.
On the other hand, for SPS with high enough values of $P_{2}$, there is a region where BB84+HPP is advantageous over WCS+D, as shown in the right corner of Fig. 3(b) (marked as ”SPS+HPP”), where $\gamma>0\,\text{dB}$. In the implementation of the HPP protocol, most one-photon events are discarded, and the key is composed largely of two-photon emission events, as indicated in Eq. 9, where the first term is dominant since $P_{DC}\ll P_{1},\,P_{2}$. Therefore, the HPP method is most advantageous in the regime where $P_{2}$ is large, regardless of $P_{1}$. The negligible probability of two-photon events after purification ($<{10}^{-7}$) and minor contributions of the second and third terms in Eq. 9 result in an effectively pure source, differing from a perfect SPS only in brightness (through the zero-photon probability).
As indicated by the first term of Eq. 9, the probability of sending one photon is linearly dependent on $\eta_{D}$, suggesting that the minimum value of $P_{2}$ required for $\gamma>0\,\text{dB}$ is inversely proportional to Alice’s detection efficiency. Calculations yield this relation, giving the condition $P_{2}\geq 0.37/\eta_{D}$ for a balanced 50:50 BS. The values of $\gamma$ and the separation line between the WCS and HPP in Fig. 3(b) are evaluated for the realistic case $\eta_{D}=0.9$, yielding an expected enhancement over WCS when $P_{2}>0.41$.
The black dot in Fig. 3(b) represents an actual SPS device consisting of a gCQD coupled to a Bragg antenna with a plasmonic nanocone (SPS2), demonstrated in Ref. [34], again showing that existing imperfect SPS+HPP can compete with WCS+D.
IV Experimental Results
After showing that existing SPSs, combined with the two new protocols can outperform WCS in terms of MCL, here we demonstrate experimental proof-of-concept emulations of both protocols using our bare single gCQD presented in Fig. 1(c) as our imperfect SPS emulator.
To experimentally demonstrate the feasibility of the DTB scheme, we use three different pulsed excitation intensities generated by a 405 nm diode laser, marked as $S_{0}=0$, $S_{1}$, and $S_{2}$ in Fig. 1(c) (see Supp. Sec. V). Fig. 4(a) demonstrates photon emission control showing stable photon counts under each excitation intensity. This allows an easy control of the photon statistics, as shown in Fig. 4(b), which presents the $g^{(2)}(0)$ values for $S_{1}$ and $S_{2}$, consistent with the modification of photon statistics shown in Fig. 1, where $P^{(S_{1})}_{1}=0.05$, $P^{(S_{2})}_{1}=0.4$ and $P^{(S_{1})}_{2}=0.0005$, $P^{(S_{2})}_{2}=0.15$.
As shown previously, the gCQD SPS with a near unity collection efficiency can outperform WCS+D, if implemented with the DTB protocol. In Fig. 4(c), we numerically calculate the expected $\gamma$ of a gCQD SPS device in terms of $\eta_{C}$ (see Supp. Sec. VI). Remarkably, with the current measured parameters, $\gamma>0\,\text{dB}$ already for $\eta_{C}>0.3$, which is well within reach. As an example, our previously demonstrated gCQD based SPS devices [36] has $\eta_{C}>0.7$ (marked by a blue rectangle) leading to an expected $\gamma>2\,\text{dB}$, a significant improvement over WCS+D.
Moving to emulation of the HPP protocol, we show experimental results of the purification of the gCQD emission in Fig. 5 (see Supp. Sec. V), using a 50:50 BS. The second-order correlation measurements of the gCQD without (Fig. 5(a)) and with (Fig. 5(b)) the HPP post-processing protocol are presented. As can be seen, a near-zero $g^{(2)}(0)$, limited only by detector noise is achieved, competing with state-of-the-art demonstrations [32, 46]. We note that with HPP, the photon rate decreases to $\sim 0.5P_{2}\eta_{D}$, but the collection efficiency $\eta_{C}$ is not affected.
Using the results in Fig. 5(a),(b), we find that $P_{3}\sim 10^{-7}$, therefore we conclude that $P_{N>2}$ is negligible comparing to $P_{1}$ and $P_{2}$ (see Supp. Sec. III), justifying our initial assumptions.
Interestingly, the HPP efficiency can be optimized by adjusting $T$, $R$ of Alice’s BS, depending on $P_{2}$ of the source. In Fig. 5(c) we show a calculation of the optimal beam splitter (BS) transmission $T$ to reach the highest MCL, as a function of $P_{2}$, for two dark-counts probabilities $P_{DC}$. For low $P_{DC}$ the optimal transmittance is roughly $0.5$, however for higher values of $P_{DC}$, the optimal $T$ is smaller than 0.5 and decreases with increasing two-photon probability, in order to minimize the probability for a dark count at Alices’ detector simultaneously with two-photons transmission to Bob.
Lastly, in Fig. 5(d) we plot he calculated $\gamma$ for a gCQD based SPS with the above parameters, as a function of $\eta_{D}$, for a fixed $\eta_{C}=0.9$ (red line), and as a function of $\eta_{C}$ for a fixed $\eta_{D}=0.9$. Again, the blue box, representing demonstrated values of SPS devices based on gCQD coupled to nano-antennas, shows an improvement over WCS+D.
V Conclusions
As an alternative to the very challenging push for nearly ideal SPS that is required to outperform the existing WCS protocols for QKD, we proposed and analyzed two simple-to-implement protocols that can allow for even far from ideal SPS, which are currently technologically ready, to beat the state-of-the-art performance of weak coherent states with decoy protocols, achieving over $>3$ dB enhancement in terms of the secure key rate.
The protocols are based on the simple ability to control the statistical distribution of the truncated photon number basis {|$\{\ket{0},\ket{1},\ket{2}\}$ of a QD BX-X cascaded emission, by varying the excitation power. We showed that depending on the possible attainable values of $P_{1}$ and $P_{2}$, either a decoy state protocol, DTB, or an heralded purification protocol, HPP, can be employed to a non-ideal SPS.
This is a particularly attractive route for improving the performance of current QKD systems, as we have shown that even room temperature, on-chip, compact, and easily integrated SPS devices, such as those based on gCQD coupled to nano-antennas, are already well within the parameter range for superior performance over WCS with decoy states by employing either a DTB protocol or HPP. Both protocols have very simple requirements and their application is very general, thus we believe they can be employed efficiently on a vast range of sub-Poisson, quantum emitters, opening a practical and realistic way to implement novel photon sources with superior QKD performance, without the stringent requirements that hindered their practical integration into real-world QKD systems.
Our protocols, which improves QKD performance, could also enhance other quantum cryptography technologies. By improving eavesdropping detection, it could strengthen quantum secure direct communication [47], quantum secret sharing [48], and quantum secure computation [49]. These protocols offer versatile improvement across various quantum cryptographic applications.